A potter's wheel having a radius of 0.500 m and ma
Answer #1

Given:-

Radius,r=0.5m

Mass,m=45 kg

Initial angular speed,wi=50rev/mt= 5.23 rad/s                    [1rev/mt= 2*pie/60 rad/s]

Force,F=70N

a]

Moment of Inertia of disk= mr2= 45*0.52=11.25 kg m2

b]

Time,t=6s

Wf= 0 rad/s

Use the kinematic equation,w_{f}=w_{i}+\alpha t

\alpha =\frac{w_{f}-w_{i}}{t}

?=0-5.23/6= -0.872 rad/s2

c]

Net Torque= Moment of inertia* angular acceleration

\tau =I*\alpha

Torque,T= 11.25 kg m2 *-0.872 rad/s2 = -9.81Nm

d]

Normal force, Fn= mv2/r= mrw2 = 45*0.5* 5.23 =117.7N

e] Coeffieint of kinetic friction between the wheel and the rag is,\mu =\frac{v^{2}}{rg}=\frac{rw^{2}}{g}

\mu =\frac{0.5*5.23^{2}}{9.8}=\mathbf{1.4}